Can you show me how to solve 630$ minus 3%,?
Answer by stanbon(75887)

solve 630$ minus 3%,?
---
630 - 0.03*630
= (1-0.03)*630
= 0.97*630
= (97/100)630
-----
= 97*6.3
= 611.1
================
Cheers,
Stan H.